OpenStudy (anonymous):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for three different years: Number of years 1 2 3 Option 1 (amount in dollars) 1100 1210 1331 Option 2 (amount in dollars) 1100 1200 1300 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points)

OpenStudy (anonymous):

Can you explain the difference between Option 1 and option 2

OpenStudy (anonymous):

option 2 is increasing by 100

OpenStudy (anonymous):

option one increases by 110 then 111 so yeah

OpenStudy (anonymous):

Which one is a linear?

OpenStudy (anonymous):

option 2 i think

OpenStudy (anonymous):

Correct. Do you know the formula for a linear function?

OpenStudy (anonymous):

y=mx=b?

OpenStudy (anonymous):

So how do you find m?

OpenStudy (anonymous):

-b so now its y-b=mx /x so now its y-b/x=m?

OpenStudy (anonymous):

something like that

OpenStudy (anonymous):

m = slope => rise over run \[\frac{ \Delta y }{ \Delta x }\] Is your time on the x axis or y axis? Is dollars on the x axis or y axis?

OpenStudy (anonymous):

dollars= y time=x

OpenStudy (anonymous):

So what is the change in y over the change in x?

OpenStudy (anonymous):

for which one opt one or opt 2

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

We're working on option 2 which is linear. You gave the correct formula for a linear equation.